# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ Abstract.
We establish general conditions for stochastic evolution equations with locally monotone drift and degenerate additive Lévy noise in variational formulation resulting in the existence of a unique invariant measure for the associated weakly ergodic Markovian Feller semigroup. We prove improved moment estimates for the solutions and the so-called $e$-property of the semigroup. Examples include the stochastic incompressible 2D Navier-Stokes equations, shear thickening stochastic power-law fluid equations, the stochastic heat equation, as well as, stochastic semilinear equations such as the 1D stochastic Burgers equation.
@@
Additive Gaussian noise; additive Poisson noise; ergodic Markovian Feller semigroup; $e$-property; locally monotone drift stochastic evolution equation with degenerate Lévy noise; stochastic Burgers equation; stochastic incompressible 2D Navier-Stokes equations; stochastic power law fluids; stochastic semilinear equations; unique invariant probability measure
@@ 1. Introduction
In this work, we study ergodicity of the Markovian dynamics associated to a general class of stochastic partial differential equations (SPDEs) on a separable Hilbert space, that is, locally monotone drift stochastic evolution equations with additive Lévy noise, which is the sum of a spatially degenerate Wiener process with possibly infinitely many modes and a degenerate compensated Poisson process with small jumps and finite second moments, which is independent from the Wiener process. Even if the degeneracy of the noise could be seen as a drawback, it is actually this contribution’s strength and main novelty. The main reason is the absence of robust techniques for proving ergodicity for those SPDEs with degenerate noise that are not strongly dissipative and thus exponentially ergodic as in [8].
@@
For non-degenerate noise, more precisely, for noise with minimal non-degeneracy assumptions on the Fourier modes of the spatial structure, the situation is entirely different. We would like to point out that this kind of minimal type of non-degeneracy is sometimes called “degenerate” (as opposed to space-time white noise) or “mildly degenerate” in the literature. In this work, we reserve the terminology “degenerate” for noise which may be zero, finite-dimensional, or spatially regular. Following this terminology, for non-degenerate noise, the main approach derives from the notion of the strong Feller property of the semigroup, combined with irreducibility, which implies uniqueness of the invariant probability measure and therefore ergodicity of the semigroup. This technique works well for both additive and multiplicative noises [20]. It has been refined by Hairer and Mattingly toward the notion of the so-called asymptotic strong Feller property [41, 40], which they used to prove the existence and uniqueness of an invariant measure for the stochastic incompressible 2D Navier-Stokes equations with additive Gaussian forcing [40]. We note that their method yields convergence of the ergodic semigroup in total variation distance, whereas the method used by us yields so-called weak-mean ergodicity, where the time-average of the dual semigroup converges weakly in the sense of probability measures. Our result relies on the ergodicity result of Komorowski, Peszat and Szarek [46], which is a refinement of the so-called lower bound technique of Lasota and Szarek [52], where the main ingredient is the so-called $e$-property of the semigroup, a type of uniform equi-continuity on bounded Lipschitz functions, combined with a asymptotic uniform lower bound for the time-average of the semigroup, which does not require tightness, as the well-known Krylov-Bogoliubov method.
@@
For strongly dissipative monotone drift SPDEs with degenerate Gaussian noise, ergodicity has been obtained, among many other contributions, in [19, 8, 53, 20, 18, 27, 25, 17, 59, 26]. See [76] for a survey article on the topic of ergodicity and Kolmogorov operators for SPDEs. For singular drift monotone SPDEs with degenerate Gaussian noise, ergodicity has been obtained in [34, 28, 35, 73, 57, 78, 29, 10, 54]. For SPDEs with jump noise, ergodicity has for instance been discussed in [31, 66]. See [72] for a recent result on ergodicity of the stochastic Allen-Cahn equation with logarithmic potential.
@@
For the stochastic Navier-Stokes equations, ergodicity has been first obtained in [40, 60, 48, 23], where the results have been further refined in [22, 12, 9, 75, 47, 41, 70, 64, 63, 5, 49]. Up to now, questions of ergodicity for stochastic Navier-Stokes equations with non-degenerate noise are an active topic of research, see [16, 30, 42, 13, 15, 44, 66].
@@
The setup for SPDEs with locally monotone drift has been introduced by Liu and Röckner in [55]. As a unifying approach, it has sparked a lot of interest and has lead to many subsequent works, see [56, 3, 14, 36, 65, 69, 50, 1, 2, 33], among many other contributions. Examples for well-posedness covered by the approach of locally monotone drift SPDEs are the stochastic incompressible 2D Navier-Stokes equations, stochastic incompressible power law fluids, the stochastic incompressible tamed 3D Navier-Stokes equations, the stochastic 3D Leray-$\alpha$ model, as well as, quasilinear equations as the stochastic $p$-Laplace equation with non-monotone perturbation, and, with certain restrictions on the growth behavior and spatial dimension, semilinear stochastic equations as the 1D stochastic Burgers equation, and the 1D stochastic Allen-Cahn equation with double-well potential, and a large class of reaction-diffusion equations.
@@
In this work, our main aim is to provide an abstract framework for ergodicity of locally monotone SPDEs with degenerate additive Lévy noise. Existence and uniqueness of solutions to our model have been obtained by Brzeźniak, Liu and Zhu in [14]. As our main example, we prove the weak ergodicity of the stochastic incompressible 2D Navier-Stokes equations with Dirichlet boundary conditions on a bounded domain or with periodic boundary conditions on a square without any spatial non-degeneracy condition on the noise, thus including the deterministic incompressible 2D Navier-Stokes equation. The examples of stochastic semilinear equations include the stochastic Burgers equation in 1D, however, due to our conditions, only small quadratic growth perturbations of the second order differential operator are permitted, thus ruling out the stochastic Allen-Cahn equation in 1D with double-well potential. This is not a surprise, however, as generally speaking, the locally monotone framework of Liu and Röckner [55, 56] is better suited for SPDEs where the highest order term is pseudo-monotone or strongly dissipative, as the construction of solutions is based on Faedo-Galerkin approximations and weak convergence methods, whereas semilinear drift SPDEs are often constructed via fixed point arguments, as for instance in the approach of Agresti and Veraar [3, 1, 2], which uses critical exponents and a local Lipschitz condition. Ergodicity for locally monotone drift SPDEs has been obtained in [80] for non-degenerate noise.
@@
Our main result Theorem 2.1 below is proved in Section 5. The proof consists of the following steps. First, we derive certain a priori estimates for the solutions. These lead to the so-called $e$-property of the semigroup. To the best of our knowledge, the $e$-property for locally monotone SPDEs is proved here for the first time. It is worthwhile to note, that the $e$-property is trivial for monotone drift SPDEs with additive noise. The main tool used here is the highly useful stochastic Gronwall Lemma by Agresti and Veraar [3], which is a special case of the more general stochastic Gronwall lemma by Geiss [32]. The next step is to compare the paths of the solutions to the solution of the deterministic partial differential equation (deterministic PDE) with noise set to zero, which has an exponential or polynomial decay behavior due to our coercivity assumptions. The conditional coupling will be proved again with the help of the stochastic Gronwall Lemma [3] and a small ball property for the Lévy noise [7]. By this kind of coupling, we are able to avoid exponential martingales, which requires the existence of exponential moments. It leads to an asymptotic irreducibility property which yields the lower bound condition of Komorowski, Peszat and Szarek [46]. This procedure to prove ergodicity has been first followed by Es-Sarhir and von Renesse in [28], and has been applied to singular multi-valued monotone SPDEs by Gess and the second author in [34].
@@
In Section 2, we present our model and the main results of this paper. In Section 3, we shall pose and discuss the main hypotheses for our results, namely Hypotheses (A)(B)(C)(D)(E)(F), and (G). Subsequently, we discuss the necessary background from Markovian semigroups and invariant measures, together with the most important auxiliary results. In Section 4, we prove Theorem 2.2, Theorem 2.3, Theorem 2.4, and Theorem 2.5. In Section 5, we prove Theorem 2.1. We collect a useful inequality in Appendix A in order to get an explicit representation of the constants appearing in the main estimates. In Appendix B, we prove a convexity result, which is used in the proof of Proposition 5.3. In Appendix C, we prove a small ball property for symmetric Lévy processes needed for the proof of Lemma 5.8.
@@
For a metric space $\mathcal{X}$, denote the continuous and bounded real-valued functions on $\mathcal{X}$ by $C_{b}(\mathcal{X})$, equipped with the supremum norm $\|f\|_{\infty}:=\sup_{x\in\mathcal{X}}|f(x)|$. Denote the Lipschitz continuous functions from $\mathcal{X}$ to $\mathbb{R}$ by $\operatorname{Lip}(\mathcal{X})$, and denote the bounded and Lipschitz continuous functions from $\mathcal{X}$ to $\mathbb{R}$ by $\operatorname{Lip}_{b}(\mathcal{X})$. For $f\in\operatorname{Lip}(\mathcal{X})$, denote the Lipschitz constant of $f$ by $\operatorname{Lip}(f)$. $\operatorname{Lip}_{b}(\mathcal{X})$ is equipped with the norm $\|f\|_{\operatorname{Lip}_{b}}:=\operatorname{Lip}(f)+\|f\|_{\infty}$. Denote by ||$\cdot$|| ⋅ |, $\cdot$, respectively, the Euclidean norm on $\mathbb{R}^{d}$, and the Euclidean scalar product on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, respectively. The transpose of a real vector or matrix is denoted by the upper index ${}^{\textup{t}}$. The adjoint of a linear operator on a Hilbert space is denoted by the upper index . For a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, we denote by $W_{0}^{1,p}({\mathcal{O}};\mathbb{R}^{k})$, $k\in\mathbb{N}$, the closure of compactly supported smooth functions $C_{0}^{\infty}({\mathcal{O}};\mathbb{R}^{k})$ in $L^{p}({\mathcal{O}};\mathbb{R}^{k})$, with respect to the Sobolev norm $\|v\|_{1,p}:=\left(\int_{\mathcal{O}}|\nabla v|^{p}\,\mathrm{d}x\right)^{1/p}$. For a Banach space $V$, denote space of linear operators from $V$ to $V$ by $L(V)$ with operator norm $\|\cdot\|_{L(V)}$. For a separable Hilbert space $H$, denote the space of Hilbert-Schmidt operators from $H$ to $H$ by $L_{2}(H)$ with Hilbert-Schmidt norm $\|\cdot\|_{2}:=\|\cdot\|_{L_{2}(H)}$. Denote by $\ell^{1}$ and $\ell^{2}$, respectively, the spaces of absolutely summable and square summable real-valued sequences, respectively. As usually, $a\wedge b$ denotes the minimum of two real numbers $a$ and $b$, and $a\vee b$ denotes the maximum of two real numbers $a$ and $b$.

### v4

# Ergodicity and mixing for locally monotone stochastic evolution equations

@@ Abstract.
We establish general quantitative conditions for stochastic evolution equations with locally monotone drift and degenerate additive Wiener noise in variational formulation resulting in the existence of a unique invariant probability measure for the associated exponentially ergodic Markovian Feller semigroup. We prove improved moment estimates for the solutions and the $e$-property of the semigroup. Furthermore, we provide quantitative upper bounds for the Wasserstein $\varepsilon$-mixing times. Examples on possibly unbounded domains include the stochastic incompressible 2D Navier-Stokes equations, shear thickening stochastic power-law fluid equations, the stochastic heat equation, as well as, stochastic semilinear equations such as the 1D stochastic Burgers equation.
@@
Additive Gaussian noise; exponentially ergodic Markovian Feller semigroup; $e$-property; locally monotone drift stochastic evolution equation with degenerate Wiener noise; stochastic Burgers equation; stochastic incompressible 2D Navier-Stokes equations; stochastic power law fluids; stochastic semilinear equations; unique invariant probability measure; mixing times
@@ 1. Introduction
In this work, we study ergodicity and quantitative upper bounds of mixing times for the Markovian dynamics associated to a general class of stochastic partial differential equations (SPDEs) on a separable Hilbert space, that is, locally monotone drift stochastic evolution equations with spatially degenerate additive Wiener noise with possibly infinitely many modes. Even if the degeneracy of the noise could be seen as a drawback, it is actually this contribution’s strength and main novelty. The main reason is the absence of robust techniques for proving ergodicity for those SPDEs with degenerate noise that are not strongly dissipative and thus exponentially ergodic as in [7, 47].
@@
For non-degenerate noise, more precisely, for noise with minimal non-degeneracy assumptions on the Fourier modes of the spatial structure, the situation is entirely different. We would like to point out that this kind of minimal type of non-degeneracy is sometimes called “degenerate” (as opposed to space-time white noise) or “mildly degenerate” in the literature. In this work, we reserve the terminology “degenerate” for noise which may be zero, finite-dimensional, or spatially regular. Following this terminology, for non-degenerate noise, the main approach derives from the notion of the strong Feller property of the semigroup, combined with irreducibility, which implies uniqueness of the invariant probability measure and therefore ergodicity of the semigroup. This technique works well for both additive and multiplicative noises [24]. Flandoli and Maslowski used this method to obtain the ergodicity for the stochastic incompressible 2D Navier-Stokes equations with additive non-degenerate Gaussian forcing [38]. Their non-degeneracy assumption was removed by Mattingly [76] and Bricmont, Kupiainen and Lefevere [13] for large viscosity. The result has been further refined by Hairer and Mattingly toward the notion of the asymptotic strong Feller property [53, 52], which they used to prove the existence and uniqueness of an invariant measure for the stochastic 2D Navier-Stokes equations with additive Gaussian forcing that admits at least four independent Fourier modes on the two-dimensional torus [52]. We note that their method yields convergence of the ergodic semigroup in total variation distance, whereas the method used by us yield weak-mean ergodicity, where the time-average of the dual semigroup converges weakly in the sense of probability measures. Our result relies on the ergodicity result of Komorowski, Peszat and Szarek [57], which is a refinement of the lower bound technique of Lasota and Szarek [65], where the main ingredient is the $e$-property of the semigroup, a type of uniform equicontinuity on bounded Lipschitz functions, combined with a asymptotic uniform lower bound for the time-average of the semigroup, which does neither require tightness, as the well-known Krylov-Bogoliubov method [24], nor the sequential weak Feller property, as the method of Maslowski and Seidler [75]. This allows us to consider a general setup which dispenses with a compact embedding of the energy space. This enables us to consider ergodicity for SPDEs on certain unbounded domains, namely, those still satisfying a Poincaré inequality, also known as domains of finite width, where the Sobolev embedding is not necessarily compact. At the borderline, our method does not yield mixing. However, we may pose quantitative conditions which yield exponential mixing and exponential ergodicity. This corresponds to the case of large viscosity relative to the noise intensity for the 2D Navier-Stokes equations. To the best of our knowledge, the possible degeneracy of the noise and the explicit bound for the mixing times are novel.
@@
For strongly dissipative monotone drift SPDEs with degenerate Gaussian noise, ergodicity has been obtained, among many other contributions, in [23, 7, 67, 24, 22, 32, 30, 20, 74, 31, 86, 85]. See [92, 19] for survey articles on the topic of ergodicity and Kolmogorov operators for SPDEs. For singular drift monotone SPDEs with degenerate Gaussian noise, ergodicity has been obtained in [42, 33, 43, 88, 71, 97, 34, 9, 68]. For SPDEs with jump noise, ergodicity has for instance been discussed in [39, 81]. See [87] for a recent result on ergodicity of the stochastic Allen-Cahn equation with logarithmic potential.
@@
For the stochastic Navier-Stokes equations, ergodicity has been first obtained in [52, 76, 60, 28, 38, 13, 48], where the results have been further refined in [27, 14, 8, 91, 59, 53, 84, 79, 78, 6, 61, 44]. Up to now, questions of ergodicity, together with questions from turbulence and advection for stochastic Navier-Stokes equations with non-degenerate noise are an active topic of research, see [11, 45, 55, 37, 10, 18, 35, 54, 15, 17, 56, 81, 21]. See [44, 36, 19] for a survey of the results.
We shall consider additive noise SPDEs of the type
$\mathrm{d}X^{x}_{t}=A(X^{x}_{t})\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X^{x}_{0}=x,$
where the initial datum $x$ lies in a separable Hilbert space $H$, and $\{W_{t}\}_{t\geq 0}$ is a cylindrical Wiener process. The detailed assumptions can be found in Sections 2 and 3. The setup for SPDEs with locally monotone drift has been introduced by Liu and Röckner in [69]. A nonlinear drift operator $A:V\to V^{\ast}$, where $V\subset H$ is a reflexive Banach space with dual $V^{\ast}$, is called locally monotone if for every $u,v\in V$,
$\langle A(u)-A(v),u-v\rangle\leq(K+\rho(u))\|u-v\|^{2}_{H},$
@@
where $K\in\mathbb{R}$, and $\rho:V\to[0,\infty)$ is locally bounded and Borel measurable on $V$. If $\rho$ depends on both $u$ and $v$ in a nontrivial way, $A$ is called fully locally monotone. As a unifying approach, it has sparked a lot of interest and has lead to many subsequent works, see [70, 4, 16, 46, 80, 83, 63, 2, 3, 41], among many other contributions. Examples for well-posedness covered by the approach of locally monotone drift SPDEs are the stochastic incompressible 2D Navier-Stokes equations, stochastic incompressible power law fluids, the stochastic incompressible tamed 3D Navier-Stokes equations, the stochastic 3D Leray-$\alpha$ model, as well as, quasilinear equations as the stochastic $p$-Laplace equation with non-monotone perturbation, and, with certain restrictions on the growth behavior and spatial dimension, semilinear stochastic equations as the 1D stochastic Burgers equation, and the 1D stochastic Allen-Cahn equation with double-well potential, and a large class of reaction-diffusion equations.
@@
In this work, our main aim is to provide an abstract framework for ergodicity of locally monotone SPDEs with degenerate additive Wiener noise. Existence and uniqueness of solutions in our setting have been obtained by Liu and Röckner in [69] and have been extended by Röckner, Shang and Zhang in [83]. As our main example, we prove the ergodicity and mixing of the stochastic incompressible 2D Navier-Stokes equations with no-slip boundary conditions on a domain of finite width or with periodic boundary conditions on a square for large enough viscosity, precisely quantified, without any spatial non-degeneracy condition on the noise, thus including the deterministic incompressible 2D Navier-Stokes equations. The examples of stochastic semilinear equations include the stochastic Burgers equation in 1D, however, due to our conditions, only small quadratic growth perturbations of the second order differential operator are permitted, thus ruling out the stochastic Allen-Cahn equation in 1D with double-well potential. This is not a surprise, however, as generally speaking, the locally monotone framework of Liu and Röckner [69, 70] is better suited for SPDEs where the highest order term is pseudo-monotone or strongly dissipative, as the construction of solutions is based on Faedo-Galerkin approximations and weak convergence methods, whereas semilinear drift SPDEs are often constructed via fixed point arguments, as for instance in the approach of Agresti and Veraar [4, 2, 3], which uses critical exponents and a local Lipschitz condition. Ergodicity for locally monotone drift SPDEs has been obtained in [98] for non-degenerate noise.
@@
Our main result Theorem 2.1 below is proved in Section 5. The proof consists of the following steps. First, we derive certain a priori estimates for the solutions. These lead to the $e$-property of the semigroup, that is, the uniform-in-time equicontinuity of the Feller semigroup $P_{t}F(x):=\mathbb{E}[F(X_{t}^{x})]$, applied to a bounded Lipschitz function $F:H\to\mathbb{R}$. It can be viewed a coupling condition at infinity. In many cases, the $e$-property is implied by the asymptotic strong Feller property by duality arguments [93]. For monotone drift SPDEs with additive noise, the $e$-property follows trivially from Gronwall’s lemma. Similarly, by the standard Gronwall lemma, one obtains an estimate of the form
$|P_{t}F(x)-P_{t}F(y)|^{2}\leq\operatorname{Lip}(F)^{2}\;\mathbb{E}\left[\|X_{t}^{x}-X_{t}^{y}\|_{H}^{2}\right]$
$\leq$ $\operatorname{Lip}(F)^{2}\;\|x-y\|^{2}_{H}\;\mathbb{E}\left[\exp\left(Kt+\int_{0}^{t}\rho(X_{s}^{x})\,\mathrm{d}s\right)\right].$
Under suitable conditions on $K<0$ and $\rho$, the time integral can be bounded pathwise by local monotonicity and Itô’s lemma by a function of the norm of the initial datum $x$ and a (local) martingale, and some lower order terms of bounded variation. The exponential moment of the martingale can be controlled by a multiple of the exponential of the quadratic variation. A similar estimate has been proposed by Butkovsky, Kulik and Scheutzow in [18, 62] for a general setup for couplings, however for non-degenerate noise.
The next step is to compare the paths of the solutions to the solution of the deterministic partial differential equation (deterministic PDE) with noise set to zero, which has an exponential or polynomial decay behavior due to our coercivity assumptions. The conditional coupling will be proved with the help of the stochastic Gronwall Lemma [40] by Geiss and the small ball property of the noise. This coupling leads to an asymptotic irreducibility property which yields the lower bound condition of Komorowski, Peszat and Szarek [57]. This procedure to prove ergodicity has been first followed by Es-Sarhir and von Renesse in [33], and has been applied to singular multi-valued monotone SPDEs by Gess and the second author in [42].
Given a prescribed error $\varepsilon>0$, we define the $\varepsilon$-mixing time (with respect to the Wasserstein-2-distance $\mathcal{W}_{2}$ on the space of probability measures with second moments) as
(1.1) $\tau^{x}_{\textsf{mix}}(\varepsilon):=\inf\{t\geq 0:\mathcal{W}_{2}(P_{t}(x,\cdot),\mu_{\ast})\leq\varepsilon\},$
where $\mu_{\ast}$ is the unique invariant measure and $P_{t}(x,A):=P_{t}\mathbbm{1}_{A}(x)$. Under certain quantitative conditions, we obtain an explicit upper bound for the $\varepsilon$-mixing time by using our improved moment bound and the exponential convergence of the semigroup, see for instance [26, 66] for the definition of mixing times in the context of Markov chains.
@@
In Section 2, we present our model and the main results of this paper. In Section 3, we shall pose and discuss the main hypotheses for our results, namely Hypotheses (A)(B)(C)(D)(E)(F)(G)(H), and (I). Subsequently, we discuss the necessary background from Markovian semigroups and invariant measures, together with the most important auxiliary results. In Section 4, we prove Theorem 2.3, Theorem 2.4, Theorem 2.5, and Theorem 2.6. In Section 5, we prove the main ergodicity Theorem 2.1. In Section 6, we prove Theorem 2.2 and Theorem 6.2 on mixing times. In Appendix A, we prove a convexity result, which is used in the proof of Proposition 5.1.
@@
For a metric space $\mathcal{X}$, denote the continuous and bounded real-valued functions on $\mathcal{X}$ by $C_{b}(\mathcal{X})$, equipped with the supremum norm $\|f\|_{\infty}:=\sup_{x\in\mathcal{X}}|f(x)|$. Denote the Lipschitz continuous functions from $\mathcal{X}$ to $\mathbb{R}$ by $\operatorname{Lip}(\mathcal{X})$, and denote the bounded and Lipschitz continuous functions from $\mathcal{X}$ to $\mathbb{R}$ by $\operatorname{Lip}_{b}(\mathcal{X})$. For $f\in\operatorname{Lip}(\mathcal{X})$, denote the Lipschitz constant of $f$ by $\operatorname{Lip}(f)$. $\operatorname{Lip}_{b}(\mathcal{X})$ is equipped with the norm $\|f\|_{\operatorname{Lip}_{b}}:=\operatorname{Lip}(f)+\|f\|_{\infty}$. Denote by $|\cdot|$, $\cdot$, respectively, the Euclidean norm on $\mathbb{R}^{d}$, and the Euclidean scalar product on $\mathbb{R}^{d}\times\mathbb{R}^{d}$, respectively. The transpose of a real vector or matrix is denoted by the upper index $^{\textup{t}}$. The adjoint of a linear operator on a Hilbert space is denoted by the upper index . For a domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, we denote by $W_{0}^{1,p}({\mathcal{O}};\mathbb{R}^{k})$, $k\in\mathbb{N}$, the closure of compactly supported smooth functions $C_{0}^{\infty}({\mathcal{O}};\mathbb{R}^{k})$ in $L^{p}({\mathcal{O}};\mathbb{R}^{k})$, with respect to the Sobolev norm $\|v\|_{1,p}:=\left(\int_{\mathcal{O}}|\nabla v|^{p}\,\mathrm{d}x\right)^{1/p}$. For Banach spaces $V_{1},V_{2}$, denote space of linear operators from $V_{1}$ to $V_{2}$ by $L(V_{1},V_{2})$ with operator norm $\|\cdot\|_{L(V_{1},V_{2})}$. For separable Hilbert spaces $H_{1},H_{2}$, denote the space of Hilbert-Schmidt operators from $H_{1}$ to $H_{2}$ by $L_{2}(H_{1},H_{2})$ with Hilbert-Schmidt norm $\|\cdot\|_{L_{2}(H_{1},H_{2})}$. As usually, $a\wedge b$ denotes the minimum of two real numbers $a$ and $b$, and $a\vee b$ denotes the maximum of two real numbers $a$ and $b$.